OpenStudy (anonymous):

given that z is proportional to 1/y complete: y: 2 4 ? 0.25 z: 8 ? 16 ?

OpenStudy (nubeer):

y=k/y first find the value of k by subsituting the valus of z and y which are given.. can u find value of k?

OpenStudy (anonymous):

no i dont know,please just tell me the answer

OpenStudy (nubeer):

z=k/y y=2 , z=8.. just put in the equation and find value of K..

OpenStudy (anonymous):

idk how..

OpenStudy (nubeer):

 do you understand this much?